OpenStudy (anonymous):

What is the ratio 16 to 40 written as a fraction in lowest terms?

OpenStudy (anonymous):

please help will give medal

OpenStudy (harleygirl18):

=16/40=2/5 since you just have to divide the top and bottom by 8. 8 is a factor of both 16 and 40. It is a common factor or in this case the greatest common factor.

OpenStudy (anonymous):

thank you!